OpenStudy (anonymous):

Find all solutions to 2 sin(theta) = sqrt(2) on the interval 0 <= theta < 2 pi theta = Give your answers as exact values, as a list separated by commas.

OpenStudy (anonymous):

2*sin x = V2 sin x = V2/2 = sin Pi/4, and sin x = sin (Pi - Pi/4) = sin 3Pi/4 a) sin x = sin Pi/4 -> x = Pi/4 b) sin x = sin 3Pi/4 -> x = 3Pi/4 x = Pi/4, 3Pi/4